OpenStudy (anonymous):

Solve. t=p+3r/6 for r

OpenStudy (anonymous):

t = p + 3r/6 So we are looking for r. Then our first step is to rewrite this as follows. t - 3r/6 - t = p - t + 3r/6 - 3r/6 Now we have: -3r/6 = p - t That gets all the terms except r on the left. Now if we multiply everything by 6, we have: -3r = 6(p-t) Which expands to: -3r = 6p - 6t Divide through by -3 and we have: -3r/-3 = (6p-6t)/-3 r = -2p + 2t Please Like Tutor Sean on facebook. It will help me help others, and I offer free help there with a community of other college tutors. Also check out http://www.tutorsean.net for free videos on math and related subjects.